The body weighing 40 kg slides across the sloping plane 4 m long, forming a 45 ° angle with the horizon. Find the work done by the force of gravity on the body during its displacement across the inclined plane.


1
Expert's answer
2021-04-05T11:12:41-0400

We can find the work done by the force of gravity on the body from the formula:


"W_g=mgh."

We can find "h" from the geometry:


"sin\theta=\dfrac{h}{d},""h=dsin\theta."

Finally, we can calculate the work done by the force of gravity on the body:


"W_g=mgdsin\theta=40\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot4.0\ m\cdot sin45^{\circ}=1109\ J."